Prove the following: 1+4+7+...+(3n – 2) n(3n-1) 2
Answer #1

1) prove 1 + 4 +7+ .. +(31-2) (31-2) = n(30-1) 2 Basis stee; nel 113-0) 311)-2. 9 1 The result n free for n=1, Induction st→ 1+4+7 + - - +30+I)-2 2 +(ह-२ ) [E(+)-२) = [34 -1) +(3! -२)+ (3k+1) + (34-1) +(3+1) 2 1+4+7+ k (34-1) +2 (3+1). . 2_ १२ 6- +17477 + -- +(3k-2) +(3k+)) = 3k75k +2. 2 36 + k + 2k +2. - 3k( k+1) + alk+1) 2. (i<+1) (136 +2) 2. 1+ 4 + 7 + -- + (3k-2) +k+